# Uncovering Grounding IDs: How External Cues Shape Multi-Modal Binding

### v2

@@ Abstract
Large vision–language models (LVLMs) show strong performance across multimodal benchmarks but remain limited in structured reasoning and precise grounding. Recent work has demonstrated that adding simple visual structures, such as partitions and annotations, improves accuracy, yet the internal mechanisms underlying these gains remain unclear. We investigate this phenomenon and propose the concept of Grounding IDs, latent identifiers induced by external cues that bind objects to their designated partitions across modalities. Through representation analysis, we find that these identifiers emerge as robust within-partition alignment in embedding space and reduce the modality gap between image and text. Causal interventions further confirm that these identifiers mediate binding between objects and symbolic cues. We show that Grounding IDs strengthen attention between related components, which in turn improves cross-modal grounding and reduces hallucinations. Taken together, our results identify Grounding IDs as a key symbolic mechanism that explains how external cues enhance multimodal binding and offers both interpretability and practical improvements in robustness.
@@ 1 Introduction
Large vision–language models (LVLMs) such as LLaVA (Liu et al., 2023b), GPT-4V (Achiam et al., 2023), and Qwen-VL (Bai et al., 2023) have achieved strong results in multimodal tasks, including image captioning, visual question answering, and embodied tasks. Despite this progress, they remain limited in structured reasoning (Campbell et al., 2024), precise grounding, and hallucination (Liu et al., 2024). Very recent studies suggest that external visual structure can systematically improve LVLM reasoning. For instance, Rudman et al. (2025) reveal that many LVLMs struggle with basic polygon identification, while adding shape annotations substantially boosts accuracy. Notably, VISER (Izadi et al., 2025) introduces more general visual structure (e.g., grids or horizontal lines) and demonstrates consistent gains across diverse reasoning tasks. Despite compelling advancements, the interpretability and mechanism-level understanding of this phenomenon remain underexplored. Prior work has largely focused on what visual structure can enhance reasoning, but not how it helps.
@@
To bridge this critical understanding gap, we introduce Grounding IDs—symbolic multimodal identifiers induced by external cues. Our central hypothesis is that structured symbols, presented either visually (e.g., row labels, partitions) or textually (e.g., prompts instructing symbol-based scanning), lead LVLMs to internally generate identifiers that bind objects to their designated partitions. These Grounding IDs propagate through embeddings and attention, enabling consistent correspondence within and across modalities. We validate this hypothesis through several analyses. First, observational probes show that Grounding IDs shape attention, reinforcing partition-level dynamics across modalities. Alignment experiments further reveal that symbolic cues strengthen within-partition binding and reduce the modality gap. Second, causal interventions confirm that external cues become causally bound to multimodal object tokens through Grounding IDs.
@@
Beyond interpretability, we extend prior work on simple scaffolds (e.g., horizontal lines) to diverse and more effective cues such as digits and character symbols. Furthermore, our ablations demonstrate that both visual and textual cues contribute to binding, with their combination yielding the strongest improvements for cross-modal grounding. Finally, we demonstrate the practical utility of Grounding IDs by showing that enhanced cross-modal attention substantially reduces hallucinations in LVLMs. Overall, these results identify Grounding IDs as a core mechanism of partition-based binding, providing both mechanistic insight into LVLM reasoning and practical gains in robustness.
@@
Figure 1: Conceptual overview of Grounding IDs. An input image is augmented with simple visual cues (e.g., {@, #, \$}) and paired with a prompt that explicitly includes these symbols. Embeddings with the same Grounding IDs are displayed in matching colors across modalities, illustrating the binding between partitions and their corresponding textual descriptions.
@@ 2 Background and Hypotheses
Earlier studies show that adding external artifacts, combined with chain-of-thought (CoT) style prompting, can substantially enhance the reasoning abilities of vision–language models by shifting them from one-pass perception toward more systematic, system-2-like processing. Rudman et al. (2025) demonstrated that LVLMs are often “shape-blind”: their vision encoders cluster frequent shapes but fail to distinguish less common ones, leading to errors in side counting and geometric reasoning. Explicit cues, such as annotated edges, encourage more deliberate strategies and yield large accuracy gains. VISER (Izadi et al., 2025) generalizes this idea beyond polygons by introducing versatile and input-agnostic visual structure, such as horizontal lines, paired with sequential scanning prompts that promote structured reasoning. This approach reduces feature binding errors, encourages serial scene parsing, and consistently improves performance across reasoning tasks like counting and visual search. Building on this foundation, our work probes the internal circuits through which simple external artifacts shape visual reasoning in LVLMs.
Recent advances in mechanistic interpretability show that LLMs solve entity–attribute binding using Binding IDs, latent vectors that link entities with their attributes (Feng & Steinhardt, 2023). Causal mediation analyses support this finding and further show that these identifiers behave additively in representation space. Follow-up work extends this idea to vision–language models, showing that they employ similar vectors linking visual objects with textual references (Saravanan et al., 2025), with properties such as factorizability and position independence. Inspired by the notion of such abstract identifiers, we investigate the underlying circuit by which external visual structures equipped with textual cues enhance LVLMs’ reasoning.
In this work, we aim to answer a key open question: why do external cues improve reasoning in LVLMs? We focus on the observation that when images and prompts are augmented with simple cues that partition them into distinct regions, LVLMs show improved performance on diverse reasoning tasks. The models appear to internally generate abstract identifiers that bind objects to their respective partitions, thereby supporting more systematic visual scanning. Fig. 1 illustrates how these identifiers emerge both in representations of partitioned areas and in the generated textual descriptions of those partitions. We refer to these identifiers as Grounding IDs since this partition-based binding enhances multimodal grounding. In particular, this study seeks to validate the following hypotheses:
Existence: Augmenting the inputs with external cues induce Grounding IDs that propagate through embeddings and attention, establishing within-partition binding across modalities.
@@ 3 Empirical Evidence for Existence of Grounding IDs
To investigate why external cues improve performance, we first analyze internal attention patterns of LVLMs. We construct a controlled synthetic dataset where each image contains 15 unique objects, sampled from 35 shape–color combinations (7 shapes $\times$ 5 colors). The model, Qwen2.5-VL (Bai et al., 2025), is tasked with generating scene descriptions. As external cues, we add symbols (&, #, \$, @) to both the image and the prompt, and separate the image into four partitions using three horizontal lines (Fig. 2). Ablations on alternative cue designs are provided in the appendix. To analyze embeddings and attention, each token must be paired with its counterpart from the other modality and assigned to the corresponding partition. Visual tokens are grouped according to their position within the four rows. Text tokens are mapped to visual objects using regular expressions, aided by the fact that objects are unique within each image. This yields a partition pseudo-label for every visual and textual token, indicating the partition to which it belongs.
Figure 2: Illustration of attention patterns under baseline and structured inputs. (a) One dataset sample (top: baseline, bottom: structured). (b) Within-modality visual attention matrices. (c) Cross-modality attention matrices. Values are averaged over 100 samples and layers 22–27 of Qwen2.5-VL.
Figure 3: Analysis of the modality gap. (a) Cross-modal alignment across layers, averaged over 100 samples, showing that improvements emerge in layers 22–27. (b) Average alignment in layers 22–27, reported separately for four partitions. Object embeddings under structured inputs achieve higher alignment than the baseline (dashed line), and symbol embeddings achieve even stronger alignment than objects.
@@
For each token, we take the maximum attention score over all heads and then average across tokens within each partition, yielding a $4\times 4$ matrix. Fig. 2 shows both within-modality and cross-modality attention matrices for baseline and structured inputs. Structured inputs exhibit stronger diagonal dominance, with attention concentrated within partitions. The same trend appears in cross-modal attention between visual and textual tokens, where structured inputs achieve stronger cross-attention than baselines. These results indicate that external cues induce systematic partition-based binding in the model’s attention.
@@
While attention captures binding structure, embedding similarity provides a complementary view of alignment between modalities. We measure cosine similarity between L2-normalized visual and textual embeddings corresponding to correctly generated object tokens. Similarities are computed layer-wise and averaged across samples. Both baseline and structured inputs show alignment strengthening in later layers (after layer 20). Structured inputs consistently achieve higher similarity, particularly in the last four layers, as shown in Fig. 3. This confirms that external cues reduce the modality gap in VLMs (Schrodi et al., 2024) by enhancing cross-modal alignment. In particular, cosine similarity between activation patches corresponding to external symbols across modalities is higher than that of dataset objects.
Figure 4: Activation swap experiment. (a) Procedure in a case where source ($c^{\prime}$) and target ($c$) contain the same objects. Activations from the & and @ partitions of $c^{\prime}$ are patched into $c$, producing the patched context $c^{*}$. Predictions in $c^{*}$ follow the transferred bindings (gray) rather than host symbols. (b) Average log-probabilities of $c$ and $c^{*}$ over valid row–symbol–object combinations. Rows and columns indicate the two selected query symbols and their corresponding objects.
@@ 4 Causal Evidence of Grounding IDs
The alignment and attention analyses in Section 3 provide correlational evidence that external cues induce partition-based alignment. Our hypothesis is that Grounding IDs act as abstract vectors that are induced to related tokens across modalities, increasing alignment and grounding. We now seek causal validation. For causal mediation analysis, we use a simplified synthetic dataset in which each image contains four rows with one object per row. Rows are labeled with non-ordinal symbols (&, #, \$, @) to prevent sequential order cues. The task is discriminative visual question answering: the prompt asks, for example, ”What is in row @?”, and the model must respond ”red circle.” To formalize notation, let us define $\textbf{o}^{s_{j}}_{s_{i}}$ as an object o that is located in the partition associated with symbol $s_{i}$ and bound to symbol $s_{j}$. For example, $\textbf{o}^{\mathdollar}_{\&}$ denotes an object positioned in the & partition but bound to \$. For convenience, we use two shorthand cases: (1) $\textbf{o}^{\sim s}_{s}$ for an object located in the partition of $s$ but bound to a different symbol, and (2) $\textbf{o}^{s}_{\sim s}$ for an object bound to $s$ but located in a different partition.
@@ 4.1 Activation Swapping Experiment
We follow a causal mediation framework similar to prior studies (Vig et al., 2020; Feng & Steinhardt, 2023). Two images $c$ and $c^{\prime}$ are randomly sampled, and their activations are extracted across all layers. We then select two random symbols (e.g., \$ and #) and swap their corresponding object patch activations between the two contexts. For instance, the object in row \$ of $c^{\prime}$ is inserted into row # of $c$, while the object in row # of $c^{\prime}$ is patched into row \$. We pass the patched input $c^{*}$ to the model and record its predictions for the selected symbols. Fig. 4(a) shows a simple case where the swapped objects are the same between $c$ and $c^{\prime}$. We intentionally choose the swap strategy to avoid ambiguity arising from the coexistence of $\textbf{o}^{s}_{\sim s}$ and $\textbf{o}^{s}_{s}$.
@@
We find that predictions consistently follow the binding symbols of the swapped objects rather than the symbols present locally in the host context. In other words, for a given partition $s$, the model outputs $\textbf{o}^{s}_{\sim s}$ instead of $\textbf{o}^{\sim s}_{s}$. For example, if the object in row @ of $c^{\prime}$ is a green square and it is inserted into row & of $c$, the model reports a green square for row @ even though that object is located beside &. This indicates that the model disregards local cues and follows the transferred symbol–object binding (Fig. 4(a)).
@@
To quantify this effect across the dataset, we consider all valid swaps (across possible objects, symbols, and locations) and measure average accuracy. If we follow the standard setting without intervention and treat $\textbf{o}^{(\cdot)}_{s}$ as the correct label, accuracy drops sharply from $\mathbf{1.00}$ to $\mathbf{0.02}$ after swapping. In contrast, if we treat $\textbf{o}^{s}_{\sim s}$ as the correct label (i.e., prediction based on transferred bindings), accuracy remains high at $\mathbf{0.98}$, confirming that predictions follow Grounding IDs; thus, object–symbol bindings are causally mediated by these IDs. Fig. 4(b) shows the average log-probability of predicted objects before and after activation swapping.
@@
Figure 5: Causal mediation analysis across layers. (a) Average logit differences between $\textbf{o}^{s}_{\sim s}$ and $\textbf{o}^{\sim s}_{s}$ across layers. (b) Layer- and head-wise signal-to-noise of attention differences between $\textbf{o}^{s}_{\sim s}$ and $\textbf{o}^{\sim s}_{s}$ patches.

### v3

# Uncovering Grounding IDs: How External Cues Shape Multimodal Binding

@@ Abstract
Large vision–language models (LVLMs) show strong performance across multimodal benchmarks but remain limited in structured reasoning and precise grounding. Recent work has demonstrated that adding simple visual structures, such as partitions and annotations, improves accuracy, yet the internal mechanisms underlying these gains remain unclear. We investigate this phenomenon and propose the concept of Grounding IDs, latent identifiers induced by external cues that bind objects to their designated partitions across modalities. Through representation analysis, we find that these identifiers emerge as consistent within-partition alignment in embedding space and reduce the modality gap between image and text. Causal interventions further confirm that these identifiers mediate binding between objects and symbolic cues. We show that Grounding IDs strengthen attention between related components, which in turn improves cross-modal grounding and reduces hallucinations. Taken together, our results identify Grounding IDs as a key symbolic mechanism that explains how external cues enhance multimodal binding and offer both interpretability and practical improvements.
@@ 1 Introduction
Large vision–language models (LVLMs), such as LLaVA (llava), GPT-4V (gpt4), and Qwen-VL (qwen_vl) have achieved strong results in multimodal tasks, including image captioning, visual question answering, and embodied tasks. Despite this progress, they remain limited in structured reasoning (binding_problem), precise grounding, and hallucination (hallucination_survey). Very recent studies suggest that external visual structure can systematically improve LVLM reasoning. For instance, forgotten_polygons reveal that many LVLMs struggle with basic polygon identification, while adding shape annotations substantially boosts accuracy. Notably, VISER (viser) introduces more general visual structure (e.g., grids or horizontal lines) and demonstrates consistent gains across diverse reasoning tasks. Despite compelling advancements, the interpretability and mechanism-level understanding of this phenomenon remain underexplored. Prior work has largely focused on what visual structure can enhance reasoning, but not how it helps.
@@
To bridge this critical understanding gap, we introduce Grounding IDs, symbolic multimodal identifiers induced by external cues. Our central hypothesis is that providing LVLMs with external visual cues (e.g., row labels, symbols, grid lines) aligned with textual cues (e.g., prompts for symbol-based scanning) leads LVLMs to generate latent identifiers that bind objects to their corresponding cues. These Grounding IDs propagate through embeddings and attention, enabling consistent correspondence within and across modalities. We validate this hypothesis through several analyses. First, observational probes show that Grounding IDs shape attention, reinforcing partition-level dynamics across modalities. Alignment experiments further reveal that symbolic cues strengthen within-partition binding and reduce the modality gap. Second, causal interventions confirm that external cues become causally bound to multimodal object tokens through Grounding IDs.
@@
Beyond interpretability, we extend prior work on simple scaffolds (e.g., horizontal lines) to more effective multimodal cues, which align input modalities through a set of unique marks. Our ablation studies show that both visual and textual cues contribute to binding, with their combination yielding the greatest improvements in cross-modal grounding. Finally, we demonstrate the practical utility of Grounding IDs by showing that enhanced cross-modal attention substantially reduces hallucinations in LVLMs. Overall, our results identify Grounding IDs as a core mechanism of partition-based binding, providing both mechanistic insight into LVLM reasoning and also practical gains on visual reasoning tasks.
@@
Figure 1: Conceptual overview of Grounding IDs. An input image is augmented with simple visual cues (e.g., {@, #, \$}) and paired with a prompt that explicitly includes these symbols. Embeddings with the same Grounding IDs are displayed in matching colors across modalities, illustrating the reinforced binding between partitions and their corresponding textual descriptions.
@@ 2 Background and Hypotheses
Earlier studies show that adding external artifacts, combined with chain-of-thought (CoT) style prompting, can substantially enhance the reasoning abilities of vision–language models by shifting them from one-pass perception toward more systematic, system-2-like processing. forgotten_polygons demonstrated that LVLMs are often “shape-blind”: their vision encoders cluster frequent shapes but fail to distinguish less common ones, leading to errors in side counting and geometric reasoning. Explicit cues, such as annotated edges, encourage more deliberate strategies and yield large accuracy gains.
VISER (viser) generalizes this idea beyond polygons by introducing versatile and input-agnostic visual structure, such as horizontal lines, paired with sequential scanning prompts that promote structured reasoning. This approach reduces feature binding errors, encourages serial scene parsing, and consistently improves performance across reasoning tasks like counting and visual search. Building on this foundation, our work probes the internal circuits through which simple external artifacts shape visual reasoning in LVLMs.
Recent advances in mechanistic interpretability show that LLMs solve entity–attribute binding using Binding IDs, latent vectors that link entities with their attributes (binding_id). Causal mediation analyses support this finding and further show that these identifiers behave additively in representation space. Follow-up work extends this idea to vision–language models, showing that they employ similar vectors linking visual objects with textual references (vision_binding_id), with properties such as factorizability and position independence. These works examine binding between items and attributes in the standard setting. Inspired by these abstract identifiers, we investigate the underlying mechanism by which external visual structures, equipped with aligned textual cues, enhance LVLMs’ reasoning.
In this work, we aim to answer a key open question: why do external cues improve reasoning in LVLMs? To this end, we introduce a general setting where both images and prompts are augmented with simple shared cues (e.g., symbol characters) that partition the input into distinct regions. With these aligned multimodal cues, the models appear to generate abstract identifiers that bind objects to their respective partitions, supporting more systematic visual scanning. Fig. 1 illustrates how these identifiers emerge both in representations of partitioned areas and in the generated textual descriptions of those partitions. We refer to these identifiers as Grounding IDs, since this partition-based binding enhances multimodal grounding. In particular, this study seeks to validate the following hypotheses:
•  Existence: Augmenting inputs with multimodal external cues induces Grounding IDs that propagate through embeddings and attention, establishing within-partition binding across modalities.
@@ 3 Empirical Evidence for Existence of Grounding IDs
Figure 2: Illustration of attention patterns under baseline and structured inputs in the scene description task. (a) One dataset sample (top: baseline, bottom: structured). (b) Within-modality visual attention matrices. (c) Cross-modality attention matrices.
To investigate why external cues improve performance, we first analyze internal attention patterns of LVLMs. We focus on inference-time reasoning only, so we use a frozen 7B Qwen2.5-VL model without any fine-tuning. The tasks involve scene description and visual question answering, where the model produces detailed descriptions of the scene based on the provided input. The analysis is conducted on a synthetic dataset of 100 images. Each image contains 15 unique objects drawn from 35 shape–color combinations (7 shapes × 5 colors). To keep the analysis straightforward, each object occupies a single patch (28×28 pixels) and does not extend into adjacent patches. The baseline input follows the standard LVLM setup: the unmodified image, together with a plain scene-description prompt. The structured input augments this baseline by adding four symbols (&, #, \$, @) to the image and prompt, and by dividing the image into four horizontal partitions with three lines, as shown in Fig. 2(a). Ablations of alternative cue designs are provided in Appendix A.4.
For the representation analysis, we first determine the textual tokens that correspond to visual objects after the model generates its output. Text tokens are matched to visual objects using regular-expression (regex) pattern matching, since each object type appears only once in the image. This process assigns a partition label to each token in both modalities, allowing us to compute attention and embedding statistics in a partition-specific manner. Visual tokens are then grouped by their location within the four rows. Finally, we analyze attention patterns and embedding similarities within visual or cross-modal regions.
Figure 3: Analysis of the modality gap. (a) Cross-modal alignment across layers, showing that improvements emerge in layers 22–27. (b) Average alignment in layers 22–27, reported separately for four partitions. Object embeddings under structured inputs achieve higher alignment than the baseline (dashed line), and symbol embeddings achieve even stronger alignment than objects.
@@
For each token, we take the maximum attention score over all heads and then average across tokens within each partition (image rows), yielding a $4\times 4$ matrix. Fig. 2 shows both within-modality and cross-modality attention matrices for baseline and structured inputs. Structured inputs exhibit stronger diagonal dominance, with attention concentrated within partitions. The same trend appears in cross-modal attention between visual and textual tokens, where structured inputs achieve stronger cross-attention than baselines. These results indicate that external cues induce systematic partition-based binding that enhances the model’s attention.
@@
While attention captures binding structure, embedding similarity provides a complementary view of alignment between modalities. We measure cosine similarity between L2-normalized visual and textual embeddings corresponding to correctly generated object tokens. Similarities are computed layer-wise and averaged across samples. Both baseline and structured inputs show alignment strengthening in later layers (after layer 20). Structured inputs consistently achieve higher similarity, particularly in the last four layers, as shown in Fig. 3. This confirms that external cues reduce the modality gap in LVLMs by enhancing cross-modal alignment. In particular, cosine similarity between activation patches corresponding to external symbols across modalities is higher than that of dataset objects. modality_gap report that post-hoc removal of gap-contributing dimensions, while effective in reducing the modality gap in VLMs, distorts the embedding geometry and degrades performance. In contrast, our results show that external multimodal cues reduce the modality gap through the model’s internal dynamics without such disruptive interventions.
Figure 4: Activation swap experiment. (a) Procedure in a case where source ($c^{\prime}$) and target ($c$) contain the same objects. Activations from the & and @ partitions of $c^{\prime}$ are patched into $c$, producing the patched context $c^{*}$. Predictions in $c^{*}$ follow the transferred bindings (gray) rather than host symbols. (b) Average log probabilities of $c$ and $c^{*}$ over valid row–symbol–object combinations. Rows and columns indicate the two selected query symbols and their corresponding objects.
@@ 4 Causal Evidence of Grounding IDs
The alignment and attention analyses in Section 3 provide correlational evidence that external cues induce partition-based alignment. Our hypothesis is that Grounding IDs act as abstract vectors that are induced to related tokens across modalities, increasing alignment and grounding. We now seek causal validation using a medium-sized LVLM, Qwen2.5-VL 7B. For causal mediation analysis, we use a simplified synthetic dataset in which each image contains four rows with one object per row. Rows are labeled with non-ordinal symbols {&,\$,#,@} to prevent sequential order cues. The task is discriminative visual question answering: the prompt asks, for example, “What is in row @?”, and the model must respond with the correct object description, such as “red circle.”
To formalize notation, we define $\mathbf{o}^{s_{j}}_{s_{i}}$ as an object $\mathbf{o}$ that is located in the partition associated with symbol $s_{i}$ (subscript) and bound by the model to symbol $s_{j}$ (superscript). For example, $\mathbf{o}^{\mathdollar}_{\&}$ denotes an object positioned in the & partition but bound to \$. For convenience, we use two shorthand cases:
•  $\mathbf{o}^{\sim s}_{s}$ (adjacent object with $s$): an object located in the partition of $s$ but bound to a different symbol.
•  $\mathbf{o}^{s}_{\sim s}$ (grounded object with $s$): an object bound to $s$ but located in a different partition.
@@ 4.1 Activation Swapping Experiment
We follow a causal mediation framework similar to prior studies (vig2020investigating; binding_id). We adopt standard terminology from mechanistic interpretability to describe the causal experiments in this section (MI_Survey). We denote by $c$ an input context consisting of an image $x$ and a text prompt $p$. Given a model $M$ and layer $\ell$, let $h^{(\ell)}(c)$ be the hidden activations at that layer, including both visual patch tokens and textual tokens. In our interventions, we use three contexts: a target context $c$, a source context $c^{\prime}$, and a patched context $c^{*}$ obtained by replacing a subset of activations in $h^{(\ell)}(c)$ with those from $h^{(\ell)}(c^{\prime})$ and then running the remaining layers of $M$ on the modified activations.
@@
Two contexts, $c$ and $c^{\prime}$ are randomly sampled, and their activations are extracted across all layers. We then select two random symbols (e.g., \$ and #) and swap the patch activations of their corresponding object between the two contexts. For instance, the object in row \$ of $c^{\prime}$ is inserted into row # of $c$, while the object in row # of $c^{\prime}$ is patched into row \$. We pass the patched input $c^{*}$ to the model and record its predictions for the selected symbols. Fig. 4(a) shows a simple case where the swapped objects are the same between $c$ and $c^{\prime}$. We intentionally choose the swap strategy to avoid ambiguity arising from the coexistence of $\textbf{o}^{s}_{\sim s}$ and $\textbf{o}^{s}_{s}$.
@@
By analyzing the predictions of the patched context $c^{*}$, we discover that predictions consistently follow the binding symbols of the swapped objects rather than the symbols present locally in the host context. In other words, for a given partition $s$, the model outputs $\textbf{o}^{s}_{\sim s}$ (grounded object) instead of $\textbf{o}^{\sim s}_{s}$ (adjacent object). For example, if the object in row @ of $c^{\prime}$ is a green square and it is inserted into row & of $c$, the model reports a green square for row @ even though that object is physically located beside &. This indicates that the model disregards local cues and follows the transferred symbol–object binding for its final prediction (Fig. 4(a)).
To quantify this effect across the dataset, we consider all valid swaps (across possible objects, symbols, and locations) and measure average swap accuracy. For each valid swap of object patches between $c$ and $c^{\prime}$, we evaluate the model on $c^{*}$ using a fixed set of query symbols $s\in\mathcal{S}$. If the model’s answer for symbol $s$ in $c^{*}$ matches the grounded object $\mathbf{o}^{s}_{\sim s}$ (i.e., it follows the transferred binding rather than the locally present object $\mathbf{o}^{\sim s}_{s}$), we count the trial as correct. Swap accuracy is the fraction of such queries for which the model outputs $\mathbf{o}^{s}_{\sim s}$.
@@
If we follow the standard setting without intervention and treat $\textbf{o}^{(\cdot)}_{s}$ as the correct label, standard accuracy drops sharply from $\mathbf{1.00}$ to $\mathbf{0.02}$ after swapping. In contrast, if we treat $\textbf{o}^{s}_{\sim s}$ as the correct label (i.e., prediction based on transferred bindings), swap accuracy remains high at $\mathbf{0.98}$, confirming that predictions follow Grounding IDs; thus, object–symbol bindings are causally mediated by these IDs. Fig. 4(b) shows the average log-probability of predicted objects before and after activation swapping.
@@
Figure 5: Causal mediation analysis across layers. (a) Average logit differences between $\textbf{o}^{s}_{\sim s}$ and $\textbf{o}^{\sim s}_{s}$ across layers, showing where the model begins to favor the grounded object. (b) Signal-to-noise scores for attention differences between $\textbf{o}^{s}_{\sim s}$ and $\textbf{o}^{\sim s}_{s}$ patches across heads and layers.
4.2 Layerwise Analysis of Grounding ID Emergence
Layerwise logit difference. To locate where in the network the swapped grounding takes effect, we apply a logit-lens probe to the response token at each layer $\ell$. Let $L^{(\ell)}(x\mid c^{*})$ be the unnormalized logit for token $x$ decoded from $h^{(\ell)}(c^{*})$. We define the layerwise logit difference
$\Delta L^{(\ell)}\;=\;L^{(\ell)}\!\bigl(\mathbf{o}^{s}_{\sim s}\mid c^{*}\bigr)\;-\;L^{(\ell)}\!\bigl(\mathbf{o}^{\sim s}_{s}\mid c^{*}\bigr),$
averaged over all valid symbol–object pairs. Positive values of $\Delta L^{(\ell)}$ indicate that, at layer $\ell$, the representation favors the grounded object over the local object.
Next, we investigate in which layers the model begins to predict the intervened grounded object ($\textbf{o}^{s}_{\sim s}$) rather than the adjacent object ($\textbf{o}^{\sim s}_{s}$). The goal is to identify the layer at which the model’s prediction shifts toward the object carrying the transferred binding in the patched context. In this experiment, we use a monochrome dataset for simplicity. We apply the logit lens technique (nostalgebraist20_logit_lens) to the single-word response token at each layer $\ell$. Let $L^{(\ell)}(x\mid c^{*})$ be the unnormalized logit for token $x$ decoded from $h^{(\ell)}(c^{*})$. We define the layerwise logit difference
$\Delta L^{(\ell)}\;=\;L^{(\ell)}\!\bigl(\mathbf{o}^{s}_{\sim s}\mid c^{*}\bigr)\;-\;L^{(\ell)}\!\bigl(\mathbf{o}^{\sim s}_{s}\mid c^{*}\bigr),$ (1)
to compute the prediction tendency between $\textbf{o}^{s}_{\sim s}$ and $\textbf{o}^{\sim s}_{s}$ in each patched context, and then average across all valid symbol–object pairs.
As shown in Fig. 5(a), this difference becomes positive in the later layers (20–27), indicating that the model increasingly favors the grounded object after the intervention. This result complements the representational trend in Fig. 3a, where alignment for structured inputs increases in the same higher layers, showing that the late-layer alignment shifts are accompanied by causal evidence of Grounding ID usage.